# Secure Multi-Key Homomorphic Encryption with Application to Privacy-Preserving Federated Learning

### v1

@@ II Related Work
A follow-up study by Chen et al. [19] introduced a multi-key TFHE and presented the implementation results. Meanwhile, another line of research [20, 21] has focused on designing multi-key variants of batch HEs, such as BFV [2, 3] and CKKS [7]. However, these constructions exhibit significant security vulnerabilities and thus applications such as privacy-preserving federated learning [27] and secure distributed sparse Gaussian processes [28] that utilize the multi-key CKKS scheme fail to meet their stated security objectives.
In this paper, we propose new multi-key variants of batch HE schemes, including CKKS and BFV, by introducing a novel masking scheme for the CKKS/BFV cryptosystem. Our scheme supports the encryption of both floating-point and integer values with multiple bits, without requiring the large ciphertext expansion as in [16, 17, 18]. Furthermore, our scheme addresses the security vulnerabilities present in existing multi-key CKKS/BFV constructions [20, 21].
III Preliminaries
In this section, we introduce the background knowledge of ring learning with errors problem, multi-key homomorphic encryption, and gadget decomposition. Table I summarizes mathematical notations used in this paper.
@@ III Preliminaries
$n$ /// $N$ The number of the clients /// The RLWE dimension
[…]
@@
$||a||_{\infty}$ The $\ell^{\infty}$-norm of the coefficient vector of $a$
$\chi$ /// $U(\cdot)$ $\chi$ distributions over $R$ / Uniform distribution
@@
$a\leftarrow A$ $a$ is sampled from a set or distribution $A$
$ct$ /// $\overline{ct}$ /// $\widehat{ct}$ Fresh ciphertext/Expanded ciphertext/Masked ciphertext
$c_{j}^{i}$ /// $\bar{c}_{j}^{i}$ The $j$th component of the ciphertext $ct_{i}$ /// $\overline{ct}_{i}$
III-A Ring Learning with Errors
The Ring Learning With Errors (RLWE) assumption is based on polynomial arithmetic with coefficients in a finite field. Specifically, let $N$ be a power of two. Define $R=\mathds{Z}[x]/(x^{N}+1)$ and $R_{Q}=\mathbb{Z}_{Q}[x]/(x^{N}+1)$. Let $\chi$ be a probability distribution over $R$, and $\sigma>0$ be a real number. The RLWE assumption, defined by the parameters $(N,Q,\chi,\sigma)$, asserts that it is computationally infeasible to distinguish between two scenarios: given polynomially many samples of either $(a,b)\in R_{Q}^{2}$ or $(a,a\cdot s+e)\in R_{Q}^{2}$, where $s$ is drawn from $\chi$ and $e$ is sampled from the discrete Gaussian distribution $D_{\sigma}$ with mean 0 and standard deviation $\sigma$, the distribution of the two cases remains indistinguishable. The RLWE assumption serves as the foundation for homomorphic encryption schemes like BFV and CKKS.
III-B Multi-Key Homomorphic Encryption
@@
  • $\overline{ct}\leftarrow\texttt{MKHE.Expand}\big{(}\{pk_{1},\cdots,pk_{n}\},$ $i,ct\big{)}$. Expands the given ciphertext $ct$, encrypted under $pk_{i}$, into an expanded ciphertext $\overline{ct}$ associated with $n$ public keys $\{pk_{1},\cdots,pk_{n}\}$.
  • $\overline{ct}\leftarrow\texttt{MKHE.Eval}(\mathcal{C},\{\overline{ct}_{1},\cdots,$ $\overline{ct}_{k}\},\{pk_{id}\}_{id\in T})$. Performs homomorphic evaluation on a circuit $\mathcal{C}$ using the ciphertexts $\{\overline{ct}_{1},\cdots,$ $\overline{ct}_{k}\}$, and outputs the ciphertext $\overline{ct}$ associated with the joint public key set $\{pk_{id}\}_{id\in T}$ that corresponds to all input ciphertexts involved.
  • $\mu:=\texttt{MKHE.Decrypt}(\overline{ct};\{sk_{id}\}_{id\in T})$. Recovers the plaintext $\mu$ using the corresponding private keys $\{sk_{id}\}_{id\in T}$ of all public keys referenced in the given ciphertext $\overline{ct}$.
A unique feature of MKHE is its use of “reference sets” $\{pk_{id}\}_{id\in T}$. Each ciphertext maintains a reference to the public keys under which the data has been encrypted. Initially, a fresh ciphertext is tied to a single key. As homomorphic computations progress and involve ciphertexts encrypted under additional keys, the reference set expands. Decryption requires all secret keys corresponding to the keys in the reference set. Specifically, in collaborative scenarios, each participant partially decrypts the ciphertext with his secret key and broadcasts the partial decryption result. Then, the plaintext can be constructed by combining all participants’ partial decryption results. This distributed decryption is as follows:
  • $\nu_{id}:=\texttt{MKHE.PartDec}(\overline{ct},sk_{id})$. Partially decrypts the ciphertext $\overline{ct}$ with the secret key $sk_{id}$ and returns the partial decryption result $\nu_{id}$.
  • $\mu:=\texttt{MKHE.FullDec}(\overline{ct},\{\nu_{id}\}_{id\in T})$. Fully decrypts to obtain the plaintext $\mu$ by combining the partial decryption results $\{\nu_{id}\}_{id\in T}$ corresponding to all ciphertexts associated with the public keys referenced in the given ciphertext $\overline{ct}$.
@@
[…]
For approximate encryption schemes such as CKKS, this correctness notion is relaxed to allow for small errors, resulting in the approximate correctness condition:
[…]
@@
Formally, let $\lambda$ be the security parameter. Consider a set of users $\mathcal{I}$ ($|\mathcal{I}|\geq 2$, i.e., multi-party setting), each holding a message $\mu_{i}$ and a public/secret key pair $(pk_{i},sk_{i})\leftarrow\texttt{MKHE.KeyGen}(pp)$, where $pp\leftarrow\texttt{MKHE.Setup}(1^{\lambda})$. The parties perform the following:
  • Message encryption: $ct_{i}\leftarrow\texttt{MKHE.Encrypt}(pk_{i},\mu_{i})$;
  • Ciphertext expansion: ${\small\overline{ct}\leftarrow\texttt{MKHE.Expand}\big{(}\{pk_{i}\}_{i\in\mathcal{I}},i,ct_{i}\big{)}}$;
  • Evaluation: $\overline{ct}\leftarrow\texttt{MKHE.Eval}(\mathcal{C},\{ct_{i}\}_{i\in\mathcal{I}};\{pk_{i}\}_{i\in\mathcal{I}})$;
  • Decryption: $\nu_{i}:=\texttt{MKHE.PartDec}(\overline{ct},sk_{i})$ and $\mu:=\texttt{MKHE.FullDec}(\overline{ct},\{\nu_{i}\}_{i\in\mathcal{I}})$.
Let $\mathcal{A}$ be a real-world adversary who observes all ciphertexts, public keys, partial decryption, and the evaluation result. Then there exists a probabilistic polynomial-time simulator $\mathsf{Sim}$ such that the following distributions are computationally indistinguishable:
  • Real-world view: The adversary’s view in the real protocol execution, including the public parameters $pp$, public keys $\{pk_{i}\}$, input ciphertexts $\{ct_{i},\overline{ct}_{i}\}$, the evaluated ciphertext $\overline{ct}$, the partial decryption $\{\nu_{i}\}$, and optionally the final output $\mu$.
  • Ideal-world simulation: The simulated view produced by $\mathsf{Sim}$, given only the public parameters $pp$ and the final output $\mu=\mathcal{C}(\mu_{1},\dots,\mu_{n})$.
@@
III-C Gadget Decomposition
In lattice-based HE schemes, the accumulation of noise during homomorphic operations poses a significant challenge to the efficiency and correctness of computations. One widely used technique for mitigating this noise growth is Gadget Decomposition. By leveraging structured representations, gadget decomposition not only reduces the complexity of operations but also provides a mechanism for controlling noise in ciphertexts, enabling efficient and accurate homomorphic computations. Informally, gadget decomposition is designed to represent elements in a ring as compact, structured linear combinations of predefined basis elements.
@@ III-C Gadget Decomposition
  • Reconstruction property: A constant vector $,\bm{g}=(g_{0},g_{1},$$\cdots,g_{\tau-1})\in R_{Q}^{\tau}$ exists such that $\langle\mathcal{H}(b),\bm{g}\rangle\equiv b\pmod{Q}$.
@@
$\begin{split}\langle sk,(c_{0},c_{1})\rangle&=\langle sk,(b\boxdot\bm{\varsigma}_{0},b\boxdot\bm{\varsigma}_{1})\rangle=\langle\mathcal{H}(b),\langle sk,\bm{\Gamma}\rangle\rangle\\ &=\langle\mathcal{H}(b),\mu\cdot\bm{g}+\bm{e}\rangle=b\cdot\mu+e\ (\mathrm{mod}\,Q),\end{split}$ (3)
@@
IV Overview of Prior Work
@@
$ct_{\text{mult}}=(c_{i,j})_{i,j\in[0,n]}\in R_{Q}^{(n+1)\times(n+1)}$, $\{evk_{i}=(\mathbf{b}_{i},\mathbf{d}_{i},\mathbf{u}_{i},\mathbf{v}_{i})\}_{i\in[1,n]}$
2:$\overline{ct}_{\text{mult}}=(\bar{c}_{i})_{i\in[0,n]}\in R_{Q}^{n+1}$
3:$\bar{c}_{0}\leftarrow c_{0}c_{0}^{\prime}$
@@
IV-A Foundational Construction of CDKS
  • $ct\leftarrow\texttt{CDKS.Encrypt}(\mu,pk)$: Samples $w\leftarrow\chi$ and $e_{0},e_{1}\leftarrow D_{\sigma}$. Encrypts the given plaintext $\mu\in R$ and outputs the ciphertext $ct=w\cdot pk+(\mu+e_{0},e_{1})\pmod{Q}$. (Encrypts the given plaintext $\mu\in R_{t}$ and returns the ciphertext $ct=w\cdot pk+(\lfloor(Q/t)\cdot\mu\rceil+e_{0},e_{1})\pmod{Q}$).
  • $\overline{ct}:=\texttt{CDKS.Expand}\left(\{pk_{1},\cdots,pk_{n}\};i;ct\right)$: Expands the given ciphertext $ct$, encrypted under $pk_{i}$, into a expanded ciphertext $\overline{ct}$ associated with $n$ public keys $\{pk_{1},\cdots,pk_{n}\}$. Specifically, a ciphertext $ct=(c_{0},c_{1})$ is expanded into the ciphertext $\overline{ct}=(\bar{c}_{0},\cdots,\bar{c}_{n})\in R_{Q}^{n+1}$, where $\bar{c}_{0}=c_{0},\bar{c}_{i}=c_{1}$, and the remaining $\{\bar{c}_{j}\}_{j\in[1,n],j\neq\{0,i\}}$ are all set to 0.
  • $\overline{ct}_{\text{mult}}\leftarrow\texttt{CDKS.Mult}(\overline{ct},\overline{ct}^{\prime};\{evk_{i}\}_{i\in[1,n]})$: For the given ciphertexts $\overline{ct}=(c_{i})_{i\in[0,n]},\overline{ct}^{\prime}=(c_{i}^{\prime})_{i\in[0,n]}\in R_{Q}^{n+1}$ and their associated evaluation keys $\{evk_{i}\}_{i\in[1,n]}$, the multiplication is performed as $ct_{\text{mult}}=(c_{i}c_{j}^{\prime})_{i,j\in[0,n]}\pmod{Q}$ ($ct_{\text{mult}}=(\lfloor(t/Q)c_{i}\cdot c_{j}^{\prime}\rceil)_{i,j\in[0,n]}\pmod{Q}$). Run Algorithm 1 with $(ct_{\text{mul}},\{evk_{i}\}_{i\in[1,n]})$ and output the result $\overline{ct}_{\text{mult}}$.
  • $\mu\leftarrow\texttt{CDKS.Decrypt}(\overline{ct},\overline{sk})$: Decrypts a given ciphertext $\overline{ct}=(\bar{c}_{0},\bar{c}_{1},\cdots,\bar{c}_{n})$ using $\overline{sk}=(1,s_{1},\cdots,s_{n})$ and outputs $\mu=\langle\overline{sk},\overline{ct}\rangle\pmod{Q}$. Return $\mu$ (Return $\mu=\lfloor(t/Q)\cdot\mu\rceil$).
@@ IV-A Foundational Construction of CDKS
$\mu=\bar{c}_{0}+\sum_{i=1}^{n}\nu_{i}=\langle\overline{sk},\overline{ct}\rangle+\sum_{i=1}^{n}e_{i}\approx\langle\overline{sk},\overline{ct}\rangle\pmod{Q}.$
IV-B Security Vulnerability of CDKS
The CDKS scheme fails to provide the level of security it claims. As a result, existing MPC applications built upon CDKS, such as secure federated learning (FL) [27] and secure distributed sparse gaussian process [28], inherit CDKS’s inherent vulnerabilities and consequently fail to achieve their intended security guarantees. This limitation arises from a fundamental flaw in CDKS that allows either the server or the clients to recover plaintexts. To illustrate this issue, consider an FL system with $n$ clients and an aggregation server (as depicted in Fig. 1) secured using CDKS. The core process unfolds as follows:

### v2

@@ 2 Related Work
A follow-up study by Chen et al. [19] introduced a multi-key TFHE. Meanwhile, another line of research [20, 21] has focused on designing multi-key variants of batch HEs, such as BFV [2, 3] and CKKS [7]. However, these constructions exhibit significant security vulnerabilities and thus MPC applications such as PPFL [27] and secure distributed sparse Gaussian processes [28] that utilize the multi-key CKKS scheme fail to meet their stated security objectives.
In this paper, we propose new multi-key variants of batch HE schemes, including BFV and CKKS, by introducing a novel masking scheme for the CKKS/BFV cryptosystem. Our scheme supports the encryption of both floating-point and integer values with multiple bits, without requiring the large ciphertext expansion as in [16, 17, 18]. Furthermore, our scheme addresses the security vulnerabilities present in existing multi-key BFV/CKKS constructions [20, 21].
3 Preliminaries
In this section, we present the necessary background for this work. A summary of the mathematical notations used throughout the paper is provided in Table I.
@@ 3 Preliminaries
$n$ // NN The number of the clients // The RLWE dimension
$\mathcal{H}$ // $\lfloor\cdot\rceil$ Gadget decomposition function // The rounding function
@@
$a||_{\infty}$ The $\ell^{\infty}$-norm of the coefficient vector of $a$
$\chi$ // $U(\cdot)$ $\chi$ distributions over $R$ / Uniform distribution
@@
$\leftarrow A aa$ is sampled from a set or distribution AA
ctct // $\overline{ct}$ // $\widehat{ct}$ Fresh ciphertext/Expanded ciphertext/Masked ciphertext
$c_{j}^{i}$ // $\bar{c}_{j}^{i}$ The $j$th component of the ciphertext $ct_{i}$ // $\overline{ct}_{i}$
3.1 Ring Learning with Errors
The Ring Learning With Errors (RLWE) assumption is based on polynomial arithmetic with coefficients in a finite field. Specifically, let $N$ be a power of two. Define $R=\mathds{Z}[x]/(x^{N}+1)$ and $R_{Q}=\mathbb{Z}_{Q}[x]/(x^{N}+1)$. Let $\chi$ be a probability distribution over $R$, and $\sigma>0$ be a real number. The RLWE assumption, defined by the parameters $(N,Q,\chi,\sigma)$, asserts that it is computationally infeasible to distinguish between two scenarios: given polynomially many samples of either $(b,a)\in R_{Q}^{2}$ or $(a\cdot s+e,a)\in R_{Q}^{2}$, where $s$ is drawn from $\chi$ and $e$ is sampled from the discrete Gaussian distribution $D_{\sigma}$ with mean 0 and standard deviation $\sigma$, the distribution of the two cases remains indistinguishable. The RLWE assumption serves as the foundation for homomorphic encryption schemes like BFV and CKKS.
3.2 Multi-Key Homomorphic Encryption
@@
  • $\overline{ct}\leftarrow\texttt{MKHE.Expand}\big(\{pk_{i}\}_{i\in[1,n]},i,ct\big)$. Expands the given ciphertext $ct$, encrypted under $pk_{i}$, into an expanded ciphertext $\overline{ct}$ associated with $n$ public keys $\{pk_{i}\}_{i\in[1,n]}$.
  • $\overline{ct}\leftarrow\texttt{MKHE.Eval}(\mathcal{C};\overline{ct}_{1},\cdots,\overline{ct}_{k};\{pk_{i}\}_{i\in[1,n]})$. Performs homomorphic evaluation on the circuit $\mathcal{C}$ over the ciphertexts $\overline{ct}_{1},\cdots,\overline{ct}_{k}$ associated with the joint public key set $\{pk_{i}\}_{i\in[1,n]}$ and outputs the evaluated ciphertext $\overline{ct}$.
  • $\mu:=\texttt{MKHE.Decrypt}(\overline{ct};\{sk_{i}\}_{i\in[1,n]})$. Recovers the plaintext $\mu$ using the corresponding private keys $\{sk_{i}\}_{i\in[1,n]}$ of all public keys referenced in the given ciphertext $\overline{ct}$.
A unique feature of MKHE is its use of “reference sets” $\{pk_{i}\}_{i\in[1,n]}$. Each expanded ciphertext maintains a reference to the public keys under which the data has been encrypted. Initially, a fresh ciphertext is tied to a single key. As homomorphic computations progress and involve ciphertexts encrypted under additional keys, the reference set expands. Decryption requires all secret keys corresponding to the keys in the reference set. Specifically, in collaborative scenarios, each party partially decrypts the ciphertext with his secret key and broadcasts the partial decryption result. Then, the plaintext can be constructed by combining all parties’ partial decryption results. This distributed decryption is as follows:
  • $\nu_{i}:=\texttt{MKHE.PartDec}(\overline{ct},sk_{i})$. Partially decrypts the ciphertext $\overline{ct}$ using the secret key $sk_{i}$ and returns the partial decryption result $\nu_{i}$.
  • $\mu:=\texttt{MKHE.FullDec}(\overline{ct},\{\nu_{i}\}_{i\in[1,n]})$. Fully decrypts to obtain the plaintext $\mu$ by combining all partial decryption results $\{\nu_{i}\}_{i\in[1,n]}$ from parties in the reference set of $\overline{ct}$.
@@
Let $\overline{ct}_{1},\cdots,\overline{ct}_{k}$ be MKHE ciphertexts encrypting messages $\mu_{1},\cdots,\mu_{k}$, respectively. Denote by $\{pk_{i}\}_{i\in[1,n]}$ the reference set associated with these ciphertexts. Suppose $\overline{ct}\leftarrow$ $\texttt{MKHE.Eval}(\mathcal{C};\overline{ct}_{1},\cdots,\overline{ct}_{k};\{pk_{i}\}_{i\in[1,n]})$ is the result of evaluating a circuit $\mathcal{C}$ over these ciphertexts. Then, decryption using the corresponding secret keys $\{sk_{i}\}_{i\in[1,n]}$ correctly recovers the result with overwhelming probability:
$\texttt{MKHE.Dec}(\overline{ct};\{sk_{i}\}_{i\in[1,n]})=\mathcal{C}(\mu_{1},\cdots,\mu_{k}).$
For approximate encryption schemes such as CKKS, this correctness notion is relaxed to allow for small errors:
$\texttt{MKHE.Dec}(\overline{ct};\{sk_{i}\}_{i\in[1,n]})\approx\mathcal{C}(\mu_{1},\cdots,\mu_{k}).$
@@
Formally, let $\lambda$ be the security parameter. Consider a set of $n$ parties ($n\geq 2$), each holding a message $\mu_{i}$ and a key pair $(pk_{i},sk_{i})\leftarrow\texttt{MKHE.KeyGen}(pp)$, where $pp\leftarrow\texttt{MKHE.Setup}(1^{\lambda})$. The parties perform the following:
  • Encryption: $ct_{i}\leftarrow\texttt{MKHE.Encrypt}(pk_{i},\mu_{i})$;
  • Expansion: $\overline{ct}_{i}\leftarrow\texttt{MKHE.Expand}\big(\{pk_{i}\}_{i\in[1,n]},i,ct_{i}\big)$;
  • Evaluation: $\overline{ct}\leftarrow\texttt{MKHE.Eval}(\mathcal{C};\{\overline{ct}_{i}\}_{i\in[1,k]};\{pk_{i}\}_{i\in[1,n]})$;
  • Decryption: $\nu_{i}:=\texttt{MKHE.PartDec}(\overline{ct},sk_{i})$ and $\mu:=\texttt{MKHE.FullDec}(\overline{ct},\{\nu_{i}\}_{i\in[1,k]})$, where $2\leq k\leq n$.
Let $\mathcal{A}$ be a real-world adversary who observes all ciphertexts, public keys, partial decryption, and the evaluation result. Then there exists a probabilistic polynomial-time (PPT) simulator $\mathsf{Sim}$ such that the following distributions are computationally indistinguishable (denoted by $\approx_{c}$):
  • Real-world view: The adversary’s view in the real protocol execution, including the public parameters $pp$, public keys $\{pk_{i}\}$, input ciphertexts $\{ct_{i},\overline{ct}_{i}\}$, evaluated ciphertext $\overline{ct}$, partial decryption $\{\nu_{i}\}$, and optionally the final output $\mu$.
  • Ideal-world simulation: The simulated view produced by $\mathsf{Sim}$, given only the public parameters $pp$ and the final output $\mu=\mathcal{C}(\mu_{1},\dots,\mu_{k})$.
@@
3.3 Gadget Decomposition
In lattice-based HE schemes, the accumulation of noise during homomorphic operations poses a significant challenge to the efficiency and correctness of computations. One widely used technique for mitigating this noise growth is Gadget Decomposition. By leveraging structured representations, it provides a mechanism to control noise in ciphertexts, enabling accurate homomorphic computations. Informally, gadget decomposition is to represent elements in a ring as compact, structured linear combinations of predefined basis elements.
@@ 3.3 Gadget Decomposition
  • Reconstruction property: A constant vector $\bm{g}=(g_{0},g_{1},$ $\cdots,g_{\tau-1})\in R_{Q}^{\tau}$ exists such that $\langle\mathcal{H}(b),\bm{g}\rangle\equiv b\pmod{Q}$.
@@
$\begin{split}\langle sk,(c_{0},c_{1})\rangle&=\langle sk,(b\boxdot\bm{\varsigma}_{0},b\boxdot\bm{\varsigma}_{1})\rangle=\langle\mathcal{H}(b),\langle sk,\bm{\Gamma}\rangle\rangle\\ &=\langle\mathcal{H}(b),\mu\cdot\bm{g}+\bm{e}\rangle=b\cdot\mu+e\ (\mathrm{mod}\,Q),\end{split}$
@@
4 Overview of Prior Work
@@
4.1 Foundational Construction of CDKS
[…]
  • $\overline{ct}:=\texttt{CDKS.Expand}\left(\{pk_{i}\}_{i\in[1,n]};j;ct\right)$: Expands the given ciphertext $ct$, encrypted under $pk_{j}$, into a expanded ciphertext $\overline{ct}$ associated with $n$ public keys $\{pk_{i}\}_{i\in[1,n]}$. Specifically, a ciphertext $ct=(c_{0},c_{1})$ is expanded into the ciphertext $\overline{ct}=(\bar{c}_{0},\cdots,\bar{c}_{n})\in R_{Q}^{n+1}$, where $\bar{c}_{0}=c_{0},\bar{c}_{j}=c_{1}$, and the remaining $\{\bar{c}_{i}\}_{i\in[1,n],i\neq\{0,j\}}$ are all set to 0.
[…]
  • $\mu\leftarrow\texttt{CDKS.Decrypt}(\overline{ct},\overline{sk})$: Decrypts a given ciphertext $\overline{ct}=(\bar{c}_{0},\bar{c}_{1},\cdots,\bar{c}_{n})$ using $\overline{sk}=(1,s_{1},\cdots,s_{n})$ to obtain $\mu=\langle\overline{sk},\overline{ct}\rangle\pmod{Q}$. Return $\mu$ (Return $\mu=\lfloor(t/Q)\cdot\mu\rceil$).
@@ 4.1 Foundational Construction of CDKS
[…]
@@
$\mu={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\bar{c}_{0}+\sum_{i=1}^{n}\nu_{i}\big){\color[rgb]{1,0,0}\rceil}\approx{\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}\rangle{\color[rgb]{1,0,0}\rceil}\pmod{Q}$.
4.2 Security Vulnerability of CDKS
The CDKS scheme fails to provide the level of security it claims. As a result, existing MPC applications built upon CDKS, such as secure FL [27] and secure distributed sparse gaussian process [28], inherit CDKS’s inherent vulnerabilities and consequently fail to achieve their intended security guarantees. This limitation arises from a fundamental flaw in CDKS that allows either the server or the clients to recover plaintexts. To illustrate this issue, consider an FL system with $n$ clients and an aggregation server (as depicted in Fig. 1) secured using CDKS. The core process unfolds as follows: